# Mixture-of-Recursions: Learning Dynamic Recursive Depths for Adaptive Token-Level Computation

### v2

@@ Abstract
Abstract: Scaling language models unlocks impressive capabilities, but the accompanying computational and memory demands make both training and deployment expensive. Existing efficiency efforts typically target either parameter sharing or adaptive computation, leaving open the question of how to attain both simultaneously. We introduce Mixture-of-Recursions (MoR), a unified framework that combines the two axes of efficiency inside a single Recursive Transformer. MoR reuses a shared stack of layers across recursion steps to achieve parameter efficiency, while lightweight routers enable adaptive token-level thinking by dynamically assigning different recursion depths to individual tokens. This allows MoR to focus quadratic attention computation only among tokens still active at a given recursion depth, further improving memory access efficiency by selectively caching only their key-value pairs. Beyond these core mechanisms, we also propose a KV sharing variant that reuses KV pairs from the first recursion, specifically designed to decrease prefill latency and memory footprint. Across model scales ranging from 135M to 1.7B parameters, MoR forms a new Pareto frontier: at equal training FLOPs and smaller model sizes, it significantly lowers validation perplexity and improves few-shot accuracy, while delivering higher throughput compared with vanilla and existing recursive baselines. These gains demonstrate that MoR is an effective path towards large-model quality without incurring large-model cost.
@@ 1 Introduction
Figure 1: Overview of Mixture-of-Recursions (MoR). (Left) Each recursion step consists of a fixed stack of layers and a router that determines whether each token should pass through or exit. This recursion block corresponds to the gray box in the middle. (Middle) The full model structure, where the shared recursion step is applied up to $N_{r}$ times for each token depending on the router decision. (Right) An example routing pattern showing token-wise recursion depth, where darker cells indicate active computation through the recursion block. Below shows the number of recursion steps of each text token, shown in colors: 1, 2, and 3.
@@
Conceptually, MoR provides a pre-training framework for latent space reasoning—performing non-verbal thinking by iteratively applying a single parameter block (Hao et al., 2024; Geiping et al., 2025; Goyal et al., 2023). However, unlike approaches that deliberate on augmented continuous prompts before generation (Liu et al., 2024b; Goyal et al., 2023; Hao et al., 2024; Shen et al., 2025), MoR enables this latent thinking directly during the decoding of each token (Zelikman et al., 2024). Furthermore, routing mechanism facilitates adaptive reasoning along the model’s vertical axis [^1] While this thinking occurs along the depth axis, it is analogous to generating continuous thoughts along the horizontal sequence axis., moving beyond the uniform, fixed thinking depth common in prior work (Geiping et al., 2025; Tack et al., 2025). In essence, MoR enables models to efficiently adjust their thinking depth on a per-token basis, unifying parameter efficiency with adaptive computation.
@@ 2.1 Preliminary
We examine four parameter-sharing strategies: Cycle, Sequence, and their variants Middle-Cycle and Middle-Sequence. Table 1 summarizes two main designs, and the full list is provided in Table 5 in the Appendix. In Cycle sharing, recursion blocks are reused cyclically. For example, consider an original non-recursive model with $L{=}9$ layers and its recursive counterpart using $N_{r}{=}3$ recursions. Under the “Cycle” strategy, the layers are shared and unrolled as [$(0,1,2),(0,1,2),(0,1,2)$]. In “Sequence” sharing, each recursion block reuses the same layer consecutively before moving to the next, resulting in [$(0,0,0),(1,1,1),(2,2,2)$] for the same configuration. Both have the same effective number of layers when unrolled ($L{=}9$), but with a different order. Furthermore, the “Middle” variants preserve full-capacity parameters at the first and last layers ($\Phi_{0}$ and $\Phi_{L-1}$), while sharing weights among the intermediate layers.
@@ 2.2 Mixture-of-Recursions
To ensure coherent progression through steps, we adopt hierarchical filtering: only tokens selected at recursion step $r$ can be re-evaluated at $r{+}1$. This simulates early-exit behavior while learning from scratch. As deeper layers tend to encode increasingly abstract and sparse information (Li et al., 2022; Yang et al., 2024; Nawrot et al., 2024), this mechanism prioritizes computation for only the most demanding tokens.
@@
$\mathcal{H}_{t}^{r+1}=\begin{cases}g_{t}^{r}f(\mathcal{H}_{t}^{r},\,\Phi^{\prime})+\mathcal{H}_{t}^{1},&\text{if }r=i\\ g_{t}^{r}f(\mathcal{H}_{t}^{r},\,\Phi^{\prime}),&\text{otherwise}\\ \end{cases}$ (2.2)
@@
Table 2: Comparison of routing strategies and key-value caching strategies. (Left) Summary of two routing strategies: expert-choice and token-choice, highlighting their pros, cons, and mitigating solutions from previous works (Raposo et al., 2024; Wang et al., 2024; Zoph et al., 2022). (Right) Relative cost efficiency of caching strategies against a vanilla Transformer (normalized to 1). Here, $N_{r}$ denotes the number of recursions, and $k$ ($k<N_{\mathrm{ctx}}$) denotes the number of selected tokens per layer. KV cache memory and IO are measured across the entire model, whereas attention FLOPs are reported per layer.
@@
Recursion-wise caching cuts KV memory and IO to approximately $(N_{r}+1)/2N_{r}$ times across the entire model (when assuming capacity factors follow a sequence like $N_{r}/N_{r}$, $\cdots$, $1/N_{r}$ over $N_{r}$ recursion steps). It also reduces per-layer attention FLOPs to a factor of $(k/N_{\text{ctx}})^{2}$ of those in vanilla models, resulting in substantially improved efficiency for both training and inference phases. Meanwhile, recursive sharing can yield maximal memory savings by globally reusing context. Specifically, significant speedups can be achieved by skipping KV projection and prefill operations at shared depths (Sun et al., 2024). However, attention FLOPs only decrease by a factor of $k/N_{\text{ctx}}$, and high volume of KV IO still leads to a decoding bottleneck.
@@ 3 Experiments
Table 3: Comparison of MoR, Recursive, and Vanilla Transformers under both fixed FLOPs (16.5e18) and token (20B) settings. All models are trained on FineWeb-Edu and evaluated by validation negative log-likelihood (NLL) and few-shot accuracy. For the isoFLOP rows, the number of training tokens ($N_{tok}$) varies by model efficiency. For the fixed-token rows, we report the effective FLOPs consumed. For the model sizes, we report non-embedding parameter counts. For the KV mechanisms, we distinguish between Cache (recursion-wise caching) and Share (recursive sharing). $\text{}^{\dagger}$n recursive models, all tokens go through fixed recursion depths ($N_{r}$), instead of adaptive depths.
@@
$\text{Recursive}^{\dagger}$ M-Cyc 2 167M 16.5 20B 2.8079 31.0 37.1 66.7 52.3 40.8 27.5 42.6
@@ 3.2 IsoFLOP Analysis
As shown in Figure 3, MoR consistently outperforms recursive baselines across all model sizes and compute budgets. While it underperforms the vanilla model at the smallest model size (135M)—likely due to a recursive capacity bottleneck—this gap closes rapidly at scale. For >360M parameters, MoR not only matches but often exceeds the Vanilla Transformer, particularly under low and mid-range budgets. Overall, these results highlight that MoR is a scalable and efficient alternative to standard Transformers. It achieves strong validation performance with significantly lower parameter counts, making it a strong candidate for both pretraining and large-scale deployment. Further details are presented in Appendix C.
@@ 3.3 Inference Throughput Evaluation
We measure throughput for 360M scale-based MoR models with recursion depths of 2, 3, and 4, trained under a 16.5e18 FLOPs budget. Throughput (tokens/second) is measured based on the generation time for tokens per sample, where the number of tokens is sampled from a normal distribution with a mean of 256, starting without any input prefix. We examine two batching configurations: a fixed batch size of 32 and a (relative) maximum batch size derived by multiplying 32 by the ratio of the maximum batch sizes of vanilla and MoR models. Further details on the experimental setup are provided in Appendix D.
@@
In 4(a), across both batch settings, all MoR variants outperform the vanilla baseline, which even leverages continuous sequence-wise batching (Yu et al., 2022; Kwon et al., 2023). Increasing recursion depth leads to more tokens exiting early and a further reduction in KV cache usage. This, in turn, boosts throughput significantly (e.g., MoR-4 achieves up to a 2.06$\times$ speedup with $B=\text{Max}$). While there’s a slight performance degradation, it can be a favorable trade-off given the substantial throughput gain. These results support that the integration of the depth-wise batching paradigm with early-exiting can significantly accelerate MoR’s actual deployment throughput.
@@ 4 Ablation Studies
Figure 4: (a) Pareto frontier of inference throughput and log-likehood for MoR and Vanilla Transformer under fixed and maximum batching scenarios. Setting details are in Appendix D. (b) Negative log-likelihood (NLL) of Recursive Transformers with $N_{r}=3$ across four different parameter-sharing strategies. We pretrained the models on 10 billion tokens. The dashed red and black lines denote the full-size Vanilla Transformer and parameter-matched vanilla models (approximately one-third scales), respectively. (c) NLL performance comparison across four different architectures with KV sharing. For MoR, green (disabled) and blue (enabled) refer to recursion-wise KV caching and recursive KV sharing strategies. MoR-E and MoR-T denotes expert-choice and token-choice MoR, respectively. All models are based on 360M scale and trained on 10 billion tokens.
@@ 4.1 Parameter Sharing Strategies
As discussed in §2.1, parameter sharing is a key component of Recursive Transformers and MoR. To identify the most effective sharing configuration, we empirically compare four aforementioned strategies: Cycle, Sequence, Middle-Cycle, and Middle-Sequence. We evaluate each strategy on Recursive Transformers based on 135M and 360M model sizes. As shown in 4(b), “Middle-Cycle” consistently achieves the lowest validation loss, and its superiority is further confirmed by the detailed results in Appendix E. Based on these findings, we adopt the “Middle-Cycle” configuration for all subsequent MoR and Recursive Transformers presented in this paper.
@@ 4.2 Routing Strategies
We conduct an extensive ablation study to understand the impact of various design choices within the expert-choice and token-choice routing schemes in our MoR framework. Detailed results are summarized in Appendix F.
@@ 4.3 KV Caching Strategies
In 4(c), we first investigate the effect of KV sharing in Vanilla and Recursive Transformers. As consistent with prior works (Brandon et al., 2024; Wu and Tu, 2024; Sun et al., 2024), if we pretrain models from the scratch, KV sharing does not often compromise performance due to the greater parameter flexibility. Surprisingly, the Recursive Transformer remains relatively robust to KV sharing, despite its reduced degrees of freedom. We found evidence for this by decomposing the KV pairs at each recursion depth into their magnitude and direction. As detailed in Appendix G, depths that share parameters exhibit highly consistent magnitude patterns and high cosine similarity, providing a clear justification for why KV sharing results in only a slight performance drop.
@@ 5.2 Routing Analysis
The allocation of recursion depth reflects token semantic importance.
[…]
@@
5(b) visualizes one example of the expert-choice router output distribution at each recursion step in a MoR model with $N_{r}=3$. For each recursion step, the normalized counts of routing scores are plotted, distinguishing between tokens selected by the expert (blue) and those not selected (orange). In all steps, auxiliary loss achieves a perfect separation in router outputs, with selected tokens sharply concentrated near a routing score of 1.0 and unselected tokens clustering near 0.0. Router output distributions for the two routing strategies and their associated design choices are detailed in Appendix H.
@@ 6 Related Work
Parameter sharing provides an orthogonal path to efficiency (Dehghani et al., 2018; Lan et al., 2019; Xia et al., 2019; Jaegle et al., 2021; Takase and Kiyono, 2021; Ng and Wang, 2024; Bae et al., 2024). The Universal Transformer first showed that repeatedly applying a single block can match the representational power of deep, non-shared stacks (Dehghani et al., 2018). Looped Transformers shown to be effective, can act as programmable computers (Giannou et al., 2023), learn iterative data-fitting algorithms (Yang et al., 2023), generalize to much longer inputs on algorithmic tasks (Fan et al., 2024), and illuminate few-shot learning by mimicking multi-step optimizers (Gatmiry et al., 2024). Furthermore, Bae et al. (2024) mitigate the accuracy loss often associated with weight tying by adding low-rank adaptation (LoRA) adapters (Hu et al., 2022) in each loop, yielding Relaxed Recursive Transformers. Recent work further demonstrates that Recursive Transformers excel at latent reasoning via recurrent depth (Geiping et al., 2025). While most prior studies focus on the efficiency gains from weight tying, the recursive architecture itself offers a second level: inspired by early-exiting (Schuster et al., 2022) and compute routing (Raposo et al., 2024), one can vary the number of recursions per input (e.g., per token), allocating compute only where it is most beneficial during both training and inference.
@@
Many works have shown that dynamic compute allocation can markedly reduce the cost of training and inference, from traditional neural networks (Bengio et al., 2015; Huang et al., 2016; Teerapittayanon et al., 2016; Panda et al., 2016) to large language models (Hou et al., 2020; Elbayad et al., 2020; Fedus et al., 2022; Bae et al., 2023; Elhoushi et al., 2024; Raposo et al., 2024). Early exiting methods learn to halt processing for “easy” samples (e.g., tokens or sequences in language modeling) by skipping the remaining layers (Elbayad et al., 2020; Schuster et al., 2022; Dehghani et al., 2018; Mofakhami et al., 2024). Alternatively, early exits can be combined with speculative decoding techniques (Chen et al., 2023; Leviathan et al., 2023) during inference by leveraging lower layers for fast drafting (Bae et al., 2023; Elhoushi et al., 2024). Recently, Mixture-of-Depths (MoD) (Raposo et al., 2024) reframed adaptive depth as a routing problem: a router at each layer selects a subset of tokens to receive the full computation, while the rest bypass the layer, yielding finer-grained conditional compute. This new form of adaptive allocation is well suited to Transformer architectures and has already been extended to other modalities (Zhang et al., 2024a; Luo et al., 2024), highlighting a promising paradigm of dynamic compute at token-level granularity. MoR applies this routing idea to recursive Transformers: tokens are dynamically sent through repeated calls of a single, weight-tied block instead of through distinct layers. This shift keeps parameter count constant, allows arbitrarily deep (adaptive) compute beyond the model’s physical depth.
@@
Key–value (KV) caching stores the per-token key and value tensors produced at each layer during autoregressive decoding; reusing them eliminates quadratic-time recomputation and boosts throughput (Shazeer, 2019; Ge et al., 2023; Liu et al., 2024a; Xiao et al., 2023; Pope et al., 2022; Kang et al., 2024; Brandon et al., 2024). Unfortunately, retaining these tensors quickly saturates GPU memory, especially for long contexts and large batches (Chowdhery et al., 2023; Brandon et al., 2024). Prior work tackles this issue by quantizing KV activations to lower precision (Hooper et al., 2024; Zhang et al., 2024b), discarding entries that contribute little to the final output (Zhang et al., 2023; Liu et al., 2023a), and sharing keys and values across attention heads (Shazeer, 2019; Ainslie et al., 2023b). Brandon et al. (2024) push this idea further, allowing adjacent layers to share the same key and value tensors and achieving additional memory savings with negligible quality loss. Our Mixture-of-Recursions offer a complementary avenue: KV caches generated in early recursions can be reused in later ones, potentially reducing memory consumption even further. This provides the advantage of only needing to run the first recursion during prefill phase (Sun et al., 2024), promising significant speedups for prompt settings over 1 million tokens. Two caching strategies in MoR can be optimized based on their distinct benefits to suit various deployment settings.
@@
An emerging line of work enables LLMs to perform reasoning internally within hidden states rather than through explicit verbalization (Goyal et al., 2023; Pfau et al., 2024; Cheng and Van Durme, 2024; Tack et al., 2025; Kong et al., ). Many approaches adopt a fixed latent reasoning depth: they insert special tokens or structured prompts (e.g., a learnable “pause” token (Goyal et al., 2023) or filler punctuation (Pfau et al., 2024)) that allow the model to execute a predetermined number of hidden reasoning passes before producing an answer. Others reuse the model’s hidden states in a closed loop for a fixed number of iterations by feeding final hidden states back as input to simulate chain-of-thought (Hao et al., 2024; Shen et al., 2025; Saunshi et al., 2025). Another line of research enhances latent reasoning by augmenting hidden states with intermediate semantic signals (Zelikman et al., 2024; Tack et al., 2025). However, these methods lack the flexibility to allocate computation where it is most needed, leading to unnecessary overhead on easy inputs and insufficient reasoning on complex ones. Building upon recent findings that looping enhances model reasoning capabilities (Chen et al., 2025; Geiping et al., 2025; Saunshi et al., 2025; Zeng et al., 2025), we believe our MoR framework provides a crucial foundation for bridging adaptive compute and latent reasoning.

### v3

@@ Abstract
Abstract: Scaling language models unlocks impressive capabilities, but the accompanying computational and memory demands make both training and deployment expensive. Existing efficiency efforts typically target either parameter sharing or adaptive computation, leaving open the question of how to attain both simultaneously. We introduce Mixture-of-Recursions (MoR), a unified framework that combines the two axes of efficiency inside a single Recursive Transformer. MoR reuses a shared stack of layers across recursion steps to achieve parameter efficiency, while lightweight routers enable adaptive token-level thinking by dynamically assigning different recursion depths to individual tokens. This allows MoR to focus quadratic attention computation only among tokens still active at a given recursion depth, further improving memory access efficiency by selectively caching only their key-value pairs. Beyond these core mechanisms, we also propose a KV sharing variant that reuses KV pairs from the first recursion, specifically designed to further decrease memory footprint. Across model scales ranging from 135M to 1.7B parameters, MoR forms a new Pareto frontier: at equal training FLOPs and smaller model sizes, it significantly lowers validation perplexity and improves few-shot accuracy, while delivering higher throughput compared with vanilla and existing recursive baselines. These gains demonstrate that MoR is an effective path towards large-model quality without incurring large-model cost.
@@ 1 Introduction
Figure 1: Overview of Mixture-of-Recursions (MoR). (Left) Each recursion step consists of a fixed stack of layers and a router that determines whether each token should pass through or exit. This recursion block corresponds to the gray box in the middle. (Middle) The full model structure, where the shared recursion step is applied up to $N_{r}$ times for each token depending on the router decision. (Right) An example routing pattern showing token-wise recursion depth, where darker cells indicate active computation through the recursion block. Below shows the number of recursion steps that each subword token undergoes to predict the next token, shown in colors: 1, 2, and 3.
@@
Conceptually, MoR provides a pre-training framework for latent space reasoning—performing non-verbal thinking by iteratively applying a single parameter block (Hao et al., 2024; Geiping et al., 2025; Goyal et al., 2023). However, unlike approaches that deliberate on augmented continuous prompts before generation (Liu et al., 2024b; Goyal et al., 2023; Hao et al., 2024; Shen et al., 2025), MoR enables this latent thinking directly during the decoding of each token (Zelikman et al., 2024). Furthermore, the routing mechanism facilitates adaptive reasoning along the model’s vertical axis [^1] While this thinking occurs along the depth axis, it is analogous to generating continuous thoughts along the horizontal sequence axis., moving beyond the uniform, fixed thinking depth common in prior work (Geiping et al., 2025; Tack et al., 2025). In essence, MoR enables models to efficiently adjust their thinking depth on a per-token basis, unifying parameter efficiency with adaptive computation.
@@ 2.1 Preliminary
We examine four parameter-sharing strategies: Cycle, Sequence, and their variants Middle-Cycle and Middle-Sequence. Table 1 summarizes two main designs, and the full list is provided in Table 5 in the Appendix. In Cycle sharing, recursion blocks are reused cyclically. For example, consider an original non-recursive model with L=9L{=}9 layers and its recursive counterpart using $N_{r}{=}3$ recursions. Under the “Cycle” strategy, the layers are shared and unrolled as [$(0,1,2),(0,1,2),(0,1,2)$]. In “Sequence” sharing, each recursion block reuses the same layer consecutively before moving to the next, resulting in [$(0,0,0),(1,1,1),(2,2,2)$] for the same configuration. Both have the same effective number of layers when unrolled (L=9L{=}9), but with a different order. Furthermore, the “Middle” variants preserve full-capacity parameters at the first and last layers ($\Phi_{0}$ and $\Phi_{L-1}$), while sharing weights among the intermediate layers.
@@ 2.2 Mixture-of-Recursions
To ensure coherent progression through steps, we adopt hierarchical filtering: only tokens selected at recursion step $r$ can be re-evaluated at r+1r{+}1. This simulates early-exit behavior while learning from scratch. As deeper layers tend to encode increasingly abstract and sparse information (Li et al., 2022; Yang et al., 2024; Nawrot et al., 2024), this mechanism prioritizes computation for only the most demanding tokens.
@@
[…]
@@
Table 2: Comparison of routing strategies and key-value caching strategies. (Left) Summary of two routing strategies: expert-choice and token-choice, highlighting their pros, cons, and mitigating solutions from previous works (Raposo et al., 2024; Wang et al., 2024; Zoph et al., 2022). (Right) Relative cost efficiency of caching strategies against a vanilla Transformer (normalized to 1). Here, $N_{r}$ denotes the number of recursions, and $k$ ($k<N_{\mathrm{ctx}}$) denotes the number of selected tokens per layer. We compare only the recursion blocks, excluding the non-shared layers. KV cache memory and IO are measured across the entire model, whereas attention FLOPs are reported per layer.
@@
Recursion-wise caching cuts KV memory and IO to approximately $(N_{r}+1)/2N_{r}$ times across the entire model (when assuming capacity factors follow a sequence like $N_{r}/N_{r}$, $\cdots$, $1/N_{r}$ over $N_{r}$ recursion steps). It also reduces per-layer attention FLOPs to a factor of $(k/N_{\text{ctx}})^{2}$ of those in vanilla models, resulting in substantially improved efficiency for both training and inference phases. Meanwhile, recursive sharing can yield maximal memory savings by globally reusing context. Speedups can be further achieved by skipping KV projection and prefill at shared depths (compatible only with Cycle strategy (Sun et al., 2024)). However, attention FLOPs only decrease by $k/N_{\text{ctx}}$, and high volume of KV IO still leads to a decoding bottleneck.
@@ 3 Experiments
[…]
@@
$\text{Recursive}^{\dagger}$ - - M-Cyc 2 167M 16.5 20B 2.8079 31.0 37.1 66.7 52.3 40.8 27.5 42.6
@@ 3.2 IsoFLOP Analysis
As shown in Figure 3, MoR consistently outperforms recursive baselines across all model sizes and compute budgets. While it underperforms the vanilla model at the smallest model size (135M)—likely due to a recursive capacity bottleneck—this gap closes rapidly at scale. For >360M parameters, MoR not only matches but often exceeds the Vanilla Transformer, particularly under low and mid-range budgets. Overall, these results highlight that MoR is a scalable and efficient alternative to standard Transformers. It achieves strong validation performance with significantly lower parameter counts, making it a strong candidate for both pretraining and large‑scale deployment. Further details are presented in Appendix D.
@@ 3.3 Inference Throughput Evaluation
We measure throughput for 360M scale-based MoR models with recursion depths of 2, 3, and 4, trained under a 16.5e18 FLOPs budget. Throughput (tokens/second) is measured based on the generation time for tokens per sample, where the number of tokens is sampled from a normal distribution with a mean of 256, starting without any input prefix. We examine two batching configurations: a fixed batch size of 32 and a (relative) maximum batch size derived by multiplying 32 by the ratio of the maximum batch sizes of vanilla and MoR models. Further details on the experimental setup are provided in Appendix E.
@@
[…]
@@ 4 Ablation Studies
Figure 4: (a) Pareto frontier of inference throughput and log-likehood for MoR and Vanilla Transformer under fixed and maximum batching scenarios. Setting details are in Appendix E. (b) Negative log-likelihood (NLL) of Recursive Transformers with $N_{r}=3$ across four different parameter-sharing strategies. We pretrained the models on 10 billion tokens. The dashed red and black lines denote the full-size Vanilla Transformer and parameter-matched vanilla models (approximately one-third scales), respectively. (c) NLL performance comparison across four different architectures with KV sharing. For MoR, green (disabled) and blue (enabled) refer to recursion-wise KV caching and recursive KV sharing strategies. MoR-E and MoR-T denotes expert-choice and token-choice MoR, respectively. All models are based on 360M scale and trained on 10 billion tokens.
@@ 4.1 Parameter Sharing Strategies
As discussed in §2.1, parameter sharing is a key component of Recursive Transformers and MoR. To identify the most effective sharing configuration, we empirically compare four aforementioned strategies: Cycle, Sequence, Middle-Cycle, and Middle-Sequence. We evaluate each strategy on Recursive Transformers based on 135M and 360M model sizes. As shown in 4(b), “Middle-Cycle” consistently achieves the lowest validation loss, and its superiority is further confirmed by the detailed results in Appendix F. Based on these findings, we adopt the “Middle-Cycle” configuration for all subsequent MoR and Recursive Transformers presented in this paper.
@@ 4.2 Routing Strategies
We conduct an extensive ablation study to understand the impact of various design choices for expert-choice and token-choice routing in our MoR framework. Detailed results are summarized in Appendix G.
@@ 4.3 KV Caching Strategies
In 4(c), we first investigate the effect of KV sharing in Vanilla and Recursive Transformers. As consistent with prior works (Brandon et al., 2024; Wu and Tu, 2024; Sun et al., 2024), if we pretrain models from the scratch, KV sharing does not often compromise performance due to the greater parameter flexibility. Surprisingly, the Recursive Transformer remains relatively robust to KV sharing, despite its reduced degrees of freedom. We found evidence for this by decomposing the KV pairs at each recursion depth into their magnitude and direction. As detailed in Appendix H, depths that share parameters exhibit highly consistent magnitude patterns and high cosine similarity, providing a clear justification for why KV sharing results in only a slight performance drop.
@@ 5.2 Routing Analysis
The allocation of recursion depth reflects contextual predictability of the subsequent token.
In the Right of Figure 1, we illustrate each token’s recursion depth, which directly indicates how easily the next token can be predicted within the given context. For example, the second subword part (e.g., “-ensively”) of a word is often straightforward to predict, thus needing fewer steps. As also shown in Table 17, while the initial generation (opening) of function words like “—”, “(”, “.”, and “,” appears easy for models, predicting their closing parts or the first token immediately following their opening might be more challenging.
@@
5(b) visualizes one example of the expert-choice router output distribution at each recursion step in a MoR model with $N_{r}=3$. For each recursion step, the normalized counts of routing scores are plotted, distinguishing between tokens selected by the expert (blue) and those not selected (orange). In all steps, auxiliary loss achieves a perfect separation in router outputs, with selected tokens sharply concentrated near a routing score of 1.0 and unselected tokens clustering near 0.0. Router output distributions for the two routing strategies and their associated design choices are detailed in Appendix I.
@@ 6 Related Work
Parameter sharing provides an orthogonal path to efficiency (Dehghani et al., 2018; Lan et al., 2019; Xia et al., 2019; Jaegle et al., 2021; Takase and Kiyono, 2021; Tan et al., 2023; Ng and Wang, 2024; Bae et al., 2024). The Universal Transformer first showed that repeatedly applying a single block can match the representational power of deep, non-shared stacks (Dehghani et al., 2018). Looped Transformers shown to be effective, can act as programmable computers (Giannou et al., 2023), learn iterative data-fitting algorithms (Yang et al., 2023), generalize to much longer inputs on algorithmic tasks (Fan et al., 2024), and illuminate few-shot learning by mimicking multi-step optimizers (Gatmiry et al., 2024). Furthermore, Bae et al. (2024) mitigate the accuracy loss often associated with weight tying by adding low-rank adaptation (LoRA) adapters (Hu et al., 2022) in each loop, yielding Relaxed Recursive Transformers. Recent work further demonstrates that Recursive Transformers excel at latent reasoning via recurrent depth (Geiping et al., 2025). While most prior studies focus on the efficiency gains from weight tying, the recursive architecture itself offers a second level: inspired by early-exiting (Schuster et al., 2022) and compute routing (Raposo et al., 2024), one can vary the number of recursions per input (e.g., per token), allocating compute only where it is most beneficial during both training and inference.
@@
Many works have shown that dynamic compute allocation can markedly reduce the cost of training and inference, from traditional neural networks (Bengio et al., 2015; Huang et al., 2016; Teerapittayanon et al., 2016; Panda et al., 2016; Graves, 2016) to large language models (Hou et al., 2020; Elbayad et al., 2020; Fedus et al., 2022; Bae et al., 2023; Elhoushi et al., 2024; Raposo et al., 2024). Early exiting methods learn to halt processing for “easy” samples (e.g., tokens or sequences in language modeling) by skipping the remaining layers (Elbayad et al., 2020; Schuster et al., 2022; Dehghani et al., 2018; Mofakhami et al., 2024). Alternatively, early exits can be combined with speculative decoding techniques (Chen et al., 2023; Leviathan et al., 2023) during inference by leveraging lower layers for fast drafting (Bae et al., 2023; Elhoushi et al., 2024). Recently, Mixture-of-Depths (MoD) (Raposo et al., 2024) reframed adaptive depth as a routing problem: a router at each layer selects a subset of tokens to receive the full computation, while the rest bypass the layer, yielding finer-grained conditional compute. This new form of adaptive allocation is well suited to Transformer architectures and has already been extended to other modalities (Zhang et al., 2024a; Luo et al., 2024), highlighting a promising paradigm of dynamic compute at token-level granularity. MoR applies this routing idea to recursive Transformers: tokens are dynamically sent through repeated calls of a single, weight‑tied block instead of through distinct layers. This shift keeps parameter count constant, allows arbitrarily deep (adaptive) compute beyond the model’s physical depth.
@@
Key–value (KV) caching stores the per-token key and value tensors produced at each layer during autoregressive decoding; reusing them eliminates quadratic-time recomputation and boosts throughput (Shazeer, 2019; Ge et al., 2023; Liu et al., 2024a; Xiao et al., 2023; Pope et al., 2022; Kang et al., 2024; Brandon et al., 2024). Unfortunately, retaining these tensors quickly saturates GPU memory, especially for long contexts and large batches (Chowdhery et al., 2023; Brandon et al., 2024). Prior work tackles this issue by quantizing KV activations to lower precision (Hooper et al., 2024; Zhang et al., 2024b), discarding entries that contribute little to the final output (Zhang et al., 2023; Liu et al., 2023a), and sharing keys and values across attention heads (Shazeer, 2019; Ainslie et al., 2023b). Brandon et al. (2024) push this idea further, allowing adjacent layers to share the same key and value tensors and achieving additional memory savings with negligible quality loss. Our Mixture-of-Recursions offer a complementary avenue: KV caches generated in early recursions can be reused in later ones, potentially reducing memory consumption even further. This also provides the advantage of only needing to run the first recursion during prefill phase (Sun et al., 2024) (only with Cycle strategy), promising significant speedups for prompt settings over 1 million tokens. Two caching strategies in MoR can be optimized based on their distinct benefits to suit various deployment settings.
@@
An emerging line of work enables LLMs to perform reasoning internally within hidden states rather than through explicit verbalization (Goyal et al., 2023; Pfau et al., 2024; Cheng and Van Durme, 2024; Tack et al., 2025; Kong et al., ). Many approaches adopt a fixed latent reasoning depth: they insert special tokens or structured prompts (e.g., a learnable “pause” token (Goyal et al., 2023) or filler punctuation (Pfau et al., 2024)) that allow the model to execute a predetermined number of hidden reasoning passes before producing an answer. Others reuse the model’s hidden states in a closed loop for a fixed number of iterations by feeding final hidden states back as input to simulate chain-of-thought (Hao et al., 2024; Shen et al., 2025; Saunshi et al., 2025). Another line of research enhances latent reasoning by augmenting hidden states with intermediate semantic signals (Zelikman et al., 2024; Tack et al., 2025). However, these methods lack the flexibility to allocate computation where it is most needed, leading to unnecessary overhead on easy inputs and insufficient reasoning on complex ones. This motivates leveraging looping mechanisms for more adaptive latent reasoning (Chen et al., 2025; Geiping et al., 2025; Saunshi et al., 2025; Zeng et al., 2025).